# Analysis Of A Long Memory Circular Convolution Model

### v1

@@ Abstract.
A stochastic model, the product of a circulant matrix and a random normal vector, is shown to produce an evolutive long memory time series with a power law spectral density. The distribution of the time series, a beta location scale family of distributions, provides a connection to the unit centered spherical distribution and directional statistics. The eigenanalysis of the deterministic circulant matrix is shown to provide estimates of the discrete Fourier spectral trend, the intrinsic dimension, the probability density shape parameter of the resulting time series, the condition number of the matrix, and a principle component analysis. Examples of the R code, used as the exploratory element of the work, are given as constructive elements of the paper. The R code may be copied, pasted into an R editor, and explored.
@@ 1. Introduction
This paper adopts the $|frequency|^{-\beta}$ definition of a long memory time series, where $\beta$, the slope of the log log graph of the power spectral density, can assume any value between zero and ten. However, as all sets of quantities described in this paper are bounded in both length and magnitude, the lowest discrete Fourier spectral frequency may approach zero as the length of the model increases, but never becomes zero. The frequency approach to zero, for time series with $\beta>0$, produces an increase in the power spectral density, the distribution of variance with respect to frequency. The result is an increase of variance as the length of the time series increases. The long memory time series is non stationary with respect to variance.
This paper’s model uses a circulant convolution [3] to generate the long memory time series. The eigenvectors of the circulant matrix that generates the convolution are functions of the discrete Fourier transform basis vectors and the eigenvalues of the matrix are the coefficients of the discrete Fourier transform. It will be shown that both the intrinsic dimension, and the probability density function, of the long memory time series are functions of the number of significant low frequency spectral components of the Fourier transform of the time series.
@@ 3. The Symmetric Circulant Matrix
The matrix is a symmetric matrix, symmetrical about its top left to lower right diagonal.The matrix is self adjoint, equal to its transpose. The first row of the matrix is identical to the first column, the second row is identical to the second column, and so on. The square symmetric positive semi-definite circulant matrix used to model long memory time series in this paper has an odd number of columns and rows. The matrix is multiplied by $rn^{-0.5}$ to produce an unitary transformation.
@@ 3.1. The Top Row of The Circulant Matrix
Table 1 is a chart of the top row of the normalized matrix ”circ” and the variable ”density” in response to two values of ”beta”. The length of the model is five. The top row of ”circ” and the variable”density” are a scaled Fourier transform pair, giving insight to the response of the model over the range of ”beta”.
@@
4. Equivalence Class of Location and Scale transforms of a Distribution
Michel Loeve [4] stated in ”Probability Theory” that the type or law of a distribution function is independent of location and scale. Expressed differently, probability distributions differing only by location and or scale transformations are an equivalence class. This definition of distribution type of a long memory time series is used in this paper. Some items to remember are:
  • The dot product of two vectors A and B is equal to the product of the $L^{2}$ norms of A and B and the cosine of the angle between A and B. The dot product of A and B = $\parallel A\parallel\parallel B\parallel cos\theta$.
@@ 6. The General Beta Distribution Of Long Memory Time Series
The histograms of figure 2 are generated from the paper’s model of long memory time series with different values of the slope of the power spectral density ”beta”. They represent four standardized density functions in the beta family of distributions. The top left is the truncated normal, the top right is the Wigner semicircle, the bottom left is the uniform, and the bottom right is the arcsine distribution. These distributions are not unique to the model used in this paper, the distributions can be generated by the R language ARFIMA model of Granger and Joyeux as well as the R language model of Timmer and Konig’s 1995 model TK95. Details can be found in an earlier paper on arXiv by the author of this paper.
@@ 6.2. Popoviciu’s Inequality of Variance
with only the Bernoulli distribution having equality with 1/4.
@@ 7. Statistical Estimates Derived from the Eigenanalysis of the Symmetric Circulant Matrix
Another solution is to estimate the statistical trends directly from the eigenanalysis of the symmetric circulant matrix used to generate the time series. The eigenvectors are the orthogonal discrete Fourier basis vectors, and the eigenvalues are the coefficients of the discrete Fourier transform.The eigenvalues of a $rn\times rn$ matrix are indexed with respect to magnitude, $\lambda_{1}$ being the largest, and $\lambda_{rn}$ being the smallest. Squaring the moduli of Fourier coefficients produces an estimate of the power spectral density. The mean integral of the power spectral density is an estimate of the variance. It will be shown that trend estimates of the beta parameter $\alpha$, intrinsic dimension d, range, and the condition number of the circulant matrix can be produced from the eigenvalues.
@@ 7.1. Estimates of Intrinsic Dimension, Beta Distribution Shape Parameter, and Principle Components
Figure 3 are two log log plots of eigenvalues for a normal distribution beta = 0.001, and an arcsine distribution, beta = 10. Replacing trace(A) in equation (11) with the sum of eigenvalues, and $A||_{1}$ by the largest eigenvalue $\lambda_{1}$ yields:
@@
Equation (13) while imprecise, is useful to observe trends at the extremes of the range of ”beta”. When beta approaches zero as in the left plot of figure 3, equation(13) approaches rn, meaning all Fourier moduli are approximately equal and the distribution will approximate a normal distribution if rn is large enough. When beta is ten, the right plot of figure 3, equation (13) approaches unity, and the sum of $\lambda_{1}$ and $\lambda_{2}$ dominate the sum of all eigenvalues.The sum of $\lambda_{1}$ plus $\lambda_{2}$, are 96.8 percent of the sum of all eigenvalues, and they are the principle components of the time series. A time series with a beta of ten has an arcsine distribution, the distribution of single frequency sinusoid. The bottom right plot of figure 1, with beta = 10, is seen to be nearly sinusoidal. The plots in figure 3 represent the Fourier spectral trends of the long memory time series. The Fourier moduli of the time series are randomized by the product of ”circ” with ”epsilon”, but on average, the spectral trend is an accurate statistic with a slope of $\beta/2$. The log log plot of the squared eigenvalues has a slope equal to $\beta$, the trend of the power spectral density.
@@ 7.2. Eigenvalue Estimate of The Variance of Long Memory Time Series
In order to estimate the variance, it is necessary to remove the coefficient of the constant offset, zero frequency Fourier term $X_{1}$. The sum of Fourier trigonometric series without the zero frequency term is zero mean. Translating equation (14) to an estimate of variance using the eigenvalues of ”circ” produces the following equation:
@@ 7.3. The Condition Number of The Symmetric Matrix ”Circ” and Geometric Collapse
A symmetric matrix has a condition number $\kappa$ equal to the maximum eigenvalue divided by the minimum eigenvalue [14].
@@
7.4. The R language Program to Evaluate the Estimates of Intrinsic Dimension, Beta Shape Parameter, and Variance Derived from the Eigenvalues of ”Circ”

### v2

@@ Abstract.
A stochastic model, the product of a circulant matrix and a random normal vector, is shown to produce an evolutive long memory time series with a power law spectral density. The distribution of the time series, a beta location scale family of distributions, yields a connection to the unit centered spherical distribution and directional statistics. The eigenanalysis of the deterministic circulant matrix is demonstrated to provide estimates of the discrete Fourier spectral trend, the intrinsic dimension, the probability density shape parameter of the resulting time series, the condition number of the matrix, and a principle component analysis. Examples of the R code, used as the exploratory element of the work, are given as constructive elements of the paper. The R code may be copied, pasted into a R editor, and explored.
@@ 1. Introduction
This paper adopts the $|frequency|^{-\beta}$ definition of a long memory time series, where $\beta$, the slope of the log log graph of the power spectral density trend, can assume any value between zero and ten. However, as all sets of quantities described in this paper are bounded in both length and magnitude, the lowest discrete Fourier spectral frequency may approach zero as the length of the model increases, but never becomes zero. The frequency approach to zero, for time series with $\beta>0$, produces an increase in the power spectral density, resulting in an increase of variance as the length of the time series increases. The long memory time series is non stationary with respect to variance.
This paper’s model uses a circular convolution [3] to generate the long memory time series. The eigenvectors of the circulant matrix that generates the convolution are the discrete Fourier transform basis vectors and the eigenvalues of the matrix are the real coefficients of the discrete Fourier transform.
By the end of the paper, it will be shown that both the intrinsic dimension, and the probability density function, of the long memory time series are functions of the number of significant low frequency spectral components of the Fourier transform of the time series, and the corresponding eigenvalues of the circulant matrix.
@@ 3. The Symmetric Circulant Matrix
The matrix is symmetrical about its main diagonal.The matrix is self adjoint, equal to its transpose. The first row of the matrix is identical to the first column, the second row is identical to the second column, and so on. The square symmetric positive definite circulant matrix used to model long memory time series in this paper has an odd number of columns and rows. The matrix is multiplied by $rn^{-0.5}$ to produce an unitary transformation.
@@ 3.1. The Top Row of The Circulant Matrix
Table 1 is a chart of the top row of the normalized matrix ”circ” and the variable ”density” in response to two values of ”beta”. The length of the model is five. The top row of ”circ” and the variable”density” are a Fourier transform pair, giving insight to the response of the model over the range of ”beta”.
@@
4. Equivalence Class of Location and Scale Transformationed Distributions
Michel Loeve [4] stated in ”Probability Theory” that the type or law of a distribution function is independent of location and scale. Expressed differently, the probability distributions of random variables differing only by location and or scale transformations are an equivalence class. The terms distribution and density in this paper refer to this equivalence class. Some items to remember are:
  • The dot product of two vectors A and B is equal to the product of the Euclidean $l^{2}$ norms of A and B and the cosine of the angle between A and B. The dot product of A and B = $\parallel A\parallel\parallel B\parallel cos\theta$.
@@ 6. The General Beta Distribution Of Long Memory Time Series
The histograms of figure 2 are generated from the paper’s model of long memory time series with different values of the slope of the power spectral density ”beta”. They represent four standardized density functions in the beta family of distributions. The top left is the truncated normal, the top right is the Wigner semicircle, the bottom left is the uniform, and the bottom right is the arcsine distribution. These distributions are not unique to the model used in this paper, the distributions can be generated by the R language ARFIMA model of Granger and Joyeux as well as the R language TK95 model of Timmer and Konig. Details can be found in an earlier paper on arXiv by the author of this paper.
@@ 6.2. Popoviciu’s Inequality of Variance
with only the Bernoulli distribution having equality with 1/4. This inequality provides an upper bound of the ratio of $\sigma^{2}$$(b-a)^{2}$ .
@@ 7. Statistical Estimates Derived from the Eigenanalysis of the Symmetric Circulant Matrix
Another solution is to estimate the statistical trends directly from the eigenvalues of the symmetric circulant matrix used to generate the time series. The eigenvalues of the $rn\times rn$ matrix are indexed with respect to magnitude, $\lambda_{1}$ being the largest, and $\lambda_{rn}$ being the smallest. What follows is a demonstration of the statistical trends that may be derived from the eigenvalues.
@@ 7.1. Estimates of Intrinsic Dimension, Beta Distribution Shape Parameter, and Principle Components
Figure 3 are two log log plots of eigenvalues for a normal distribution beta = 0.001, and an arcsine distribution, beta = 10. Replacing trace(A) in equation (11) with the sum of eigenvalues, and $A||_{2}$ by the largest eigenvalue $\lambda_{1}$ yields:
@@
Equation (13) while imprecise, is useful to observe trends at the extremes of the range of ”beta”. When beta approaches zero as in the left plot of figure 3, equation(13) approaches rn, all Fourier moduli are approximately equal, and the distribution tends to normal. When beta is ten, the right plot of figure 3, equation (13) approaches unity, the sum of $\lambda_{1}$ and $\lambda_{2}$ dominate the sum of all eigenvalues, and they are the principle components of the time series. A time series with a beta of ten has an arcsine distribution, the distribution of single frequency sinusoid. The bottom right plot of figure 1, with beta = 10, is seen to be nearly sinusoidal. The plots in figure 3 represent the Fourier spectral trends of the long memory time series. The Fourier moduli of the time series are randomized by the product of ”circ” with ”epsilon”, but on average, the spectral trend is an accurate statistic with a slope of $\beta/2$. The log log plot of the squared eigenvalues has a slope equal to $\beta$, the trend of the power spectral density.
@@ 7.2. Eigenvalue Estimate of The Variance of Long Memory Time Series
In order to estimate the variance, it is necessary to remove the coefficient of the constant offset, zero frequency Fourier term $X_{1}$. Translating equation (14) to an estimate of variance using the eigenvalues of ”circ” that correspond to the Fourier coefficients $X_{k}$, produces the following equation:
@@ 7.3. The Condition Number of The Symmetric Matrix ”Circ” and Geometric Collapse
A positive definite matrix, a matrix with positive definite eigenvalues, has a condition number $\kappa$ equal to the maximum eigenvalue divided by the minimum eigenvalue [14].
@@
7.4. The R language Program to Evaluate the Estimates of Intrinsic Dimension, Beta Shape Parameter, and Variance Derived from the Eigenvalues of ”Circ”, and Measured from Sample Variance and Range Squared